AmigaWorld

\begin{matrix}c \\ ◯ \end{matrix} by Wolfgang Lug

Documentation for version 5.19

Abstract

AmigaWorld is an informative and up-to-date database with much information about every country, many descriptions and maps.

This documentation was made to show you those abilities of Amiga-World that you might not notice by experimenting. Please have a look at the example in the appendix.

Viva las Amigas!

Contents

1 Basics 4 1.1 What AmigaWorld is . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.2 What AmigaWorld is not . . . . . . . . . . . . . . . . . . . . . . . 4 1.3 System requirements . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.4 Getting started . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.5 Changing to another language . . . . . . . . . . . . . . . . . . . . 5

2 The screen 5 2.1 The country list . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.2 The search gadgets . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.3 The little world map . . . . . . . . . . . . . . . . . . . . . . . . . 6

3 The menus 6 3.1 “AmigaWorld” . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.2 “Selection” . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.3 “Sort by” . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.4 “Display” . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 3.5 “Configuration” . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

4 Country windows 9

5 Data 9 5.1 Basic data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 5.2 Population data . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 5.3 Economic data . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

6 Maps 10

7 Description Texts 10

8 Legal stuff 11 8.1 Disclaimer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 8.2 Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

9 Registration 11

10 Creating a new translation 12

11 A small request 12

12 A big “Dankeschön” 12

A Example 15 CONTENTS 3 B Keyboard commands 16 C AmigaWorld on the WorldWideWeb 16

1 Basics

1.1 What AmigaWorld is

AmigaWorld contains much information about every country on Earth that you can call up in a country window. You can easily compare data of two or more countries by opening up to ten of these windows. More advanced comparation features are offered by the country list: You can select particular countries and sort them or search through them in many different ways.

Additionaly AmigaWorld contains informative description texts about every country and about all continents and many international organizations. Maps can show you the location of a country, vice versa you can select countries by clicking on a map.

AmigaWorld is easy to use and offers output in English, German, French, Dutch, Italian, Swedish, Norwegian and Finnish. Screen mode, font and colours can be chosen by you.

The demo-version is limited: You cannot access all of AmigaWorld’s features. Please have a look at the file called “Limitations” to see what you are missing.

1.2 What AmigaWorld is not

AmigaWorld cannot take the place of an atlas because it is not the program’s goal to display sophisticated maps, but to give fast access to information and the ability to compare data that you had to look up for each country seperately.

There is a certain PC program that knows much more data but if you do not have to know how many pharmacists work in Birmingham: AmigaWorld does know the important facts.

1.3 System requirements

AmigaWorld Version 5 requires at least Workbench 2.1, 2 MByte of RAM and a hard disk. A 68000 processor will do fine.

1.4 Getting started

You cannot start AmigaWorld directly form the diskettes that you received it on: You have to install the software on your hard disk first. This is done the usual way with the Installer program. If you have received AmigaWorld on a CD-ROM or via Aminet then you can get going immediately.

Simply double-click on its icon to start AmigaWorld from Workbench. If you like to run it from Shell you have to enter cd AmigaWorld: (i.e. the complete path where the program is located) and stack 8000 before starting the program. Parameters cannot be specified.

1.5 Changing to another language

After start AmigaWorld introduces itself in English. But you can switch to another user interface! Select “Language” from the “Configuration” menu and click on one of the available languages. All data and some of the descriptions will now be displayed in the language of your choice.

Just by the way this feature of AmigaWorld offers you the great opportunity to study the basics of a foreign language! Why not try to cope with the German user interface!

2 The screen

The most important elements of the AmigaWorld screen are the country list, the search gadgets and the little world map.

2.1 The country list

All countries that AmigaWorld knows about are displayed in the list at first. Clicking on a country with the mouse brings up a country window with the basic data about it. The menu of this window offers you more information.

You can affect the contents of the list with the “Selection” menu, e.g. have a look at the countries of one continent or one organization only. The order of the list is fixed in the “Sort by” menu.

The “Display” menu determines the look of the list: You can choose if dependencies and rank numbers are to be displayed, and so on.

2.2 The search gadgets

With the help of the text gadget you can search through the selected countries for different things. You can e.g. find out where Spanish is spoken. Wheter to search for a country’s name, a language or something else is chosen in the gadget below.

Search Mode

There are two search modes: “exclude” means that AmigaWorld only searches the selected countries. Countries that do not match your input will be excluded from the list. By using this mode you can for example find out in which countries German and Italian is spoken: Select all countries, choose “Search for language”, then type “Italian” into the gadget, hit return, type “German” and hit return again.

The “include” mode works contrastingly: All countries in the list remain selected, and all other countries that match your input will be included too.

So you can find out where German or Italian is spoken: Select all countries and switch on “mode: exclude”, and search for Italian. Then switch on “mode: include” and search for German.

Pecularities

AmigaWorld knows about accents and umlauts: It does not matter if you enter ”e” or ”é”, ”ae”, ”æ” or ”ä”. Search is not case-sensitive, i.e. you can enter ”Åland” or ”aAlaND”.

If you do not want to type in the whole word: The first letters will do, like “Great B*”. Other wildcards are not supported. Some things are to be known when searching for …

Name: If you enter the abbreviation of a country instead of its name (like “GB”) then the country itself and all its dependencies will be found. AmigaWorld knows more than one name for many countries: It does not matter if you type in “England”, “Great Britain” or “United K*”.

Location: Possible input is “East Asia” or “Central Europe”, but you can also type in the name of a sea,^{1} to get a list of the adjacent countries.

Religion: Do not enter the name of a religion (“Islam”), but the name of its followers (“Muslimes”).

Currency: You can only search for main-units (“Dollar”), not for sub-units (“Cents”).

2.3 The little world map

Clicking the left mouse button over the map brings up a huge map of the continent you aimed at. Now you can call up the data of a country by clicking it. Clicking the right mouse button brings up a big world map.

3 The menus

AmigaWorld offers the following five menus:

3.1 “AmigaWorld”

This menu provides some information on AmigaWorld and the most important feature: “Quit”.

1AmigaWorld knows the following seas: Atlantic, North Sea, Baltic Sea, Mediterranean, Black Sea, Indic, Red Sea, Pacific.

3.2 “Selection”

If you would like to have a look at the countries of one continent or organization only, then you can get a list of them with this menu. The first two items were designed for use after searching.

All countries: puts all 258 countries and dependencies into the selection list.

Restore: puts exactly those countries back to the list that have been there before your last search.

Continent: brings up a new window with a list of the continents. Click one to get a description of it and a list of the countries located there.

Organization: works the same way, with a choice of important international organizations.

Calc average: This item makes it possible to have a look at the average^{2} value of any figure for the selected countries. So you can find out the average life expectancy in the OECD countries and compare that value to the one for Africa. Use the sort menu before, for specifying the figure you want to look at.

Calc sum: This item works similar, except for the fact that you can only look at the sum of area, population and GNP – adding up the other figures would be senseless, anyway.

Load / Save: Saves or loads the contents of the list.

3.3 “Sort by”

Gives you 14 different criteria to sort the list: When selecting size, Russia comes first and the Vatican is at the end. If you can’t stand the pope being displayed at the bottom then you might use the “Reverse order” item …

2It is the “real” average, not the arithmetical mean: E.g. when calculating life expectancy you would not get the right result by adding all values and dividing their sum by the number of countries. Example: Country A has 100 inhabitants with an expected lifetime of 60 years, in country B live 20 human beings (m) that become 80 years (a) old. The average is not the mean of 60 and 80, but 63.3:

\frac{100m∗60a+20m∗80a}{100m+20m} = \frac{6000am+1600am}{120m} = \frac{7200a}{120} = 63, \overline{3}a.

Don’t panic—No more mathematics in this doc file!

3.4 “Display”

This menu determines the look of the country list. It is most interesting in connection with the “Sort by” menu as the example in the appendix will show.

Rank: Each country gets a number in fromt of it when this option is active.

Official names: The official names are displayed instead of the common ones. Quite useless.

Value: When you activate this option and the list is sorted by population then the number of inhabitants will be displayed behind every country. This works the same way with all other choices that you make in the “Sort by” by, except for alphabetical order, of course.

Dependencies: By de-activating this one, you can remove all dependencies from the selection list.

Compare: Not only the currently selected countries will be displayed, but also those that had been selected before your last search or your last use of the “Selection” menu. They will be displayed in a different colour, so you can easily compare two lists.^{3} Moreover you can (de)select single countries with the right mouse button.

3.5 “Configuration”

With this one you can adjust the looks of AmigaWorld quite largely to your personal preferences:

Palette: On 16-colour screens maps and flags are still displayed in the right colours, but on an 8-coloured screen they will not look very good after changing the palette.

Screen Mode: Another requester appears, and you can choose a new screen mode. Better not use lores-modes.

Font: “FarawayEyes” font that comes along with the program is recommended.

Language: For switching to another user-interface.

Measures: A feature for all you Anglo-Americans: You can get area figures in sqare miles instead of square kilometres.

Save: Puts the settings of the config menu to disk; AmigaWorld will use them right after you start it the next time.

3If you want two compare two completely different lists that can only be created with more than one step: Save list 1 to disk, create list 2, load list 1 again, restore list 2 with the second item of the “Selection” menu—then you can compare.

4 Country windows

At first, the basic data is displayed in the country window. Using the “Data” menu gives you population or economic data instead.

The other menu offers even more information: You can call up an English description about the country, a location map and a list of its dependencies, if any. Moreover you can print out the data.

You can still use all menus of the main window, the country list and the search feature, after having opened up to ten country windows.

5 Data

Most data was taken from the 1998 edition of “Fischer-Weltalmanach”, a German yearbook. Other sources include the internet, newspaper articles and information provided by users. If you are interested to understand what kind of data is given you then you should read the following sections:

5.1 Basic data

Several countries have changed their names recently, like Burma into Myanmar or Zaire into Democratic Republic of Congo. If you cannot find a country in the list, you are probably searching for an old name. No problem: Type the name you know into the input gadget—AmigaWorld probably knows it and shows you the data of the country, along with its new name.

Not all languages spoken in a country are listed, but the most common ones. Minority languages are usually mentioned if their speakers are a local group. Not mentioned are the languages of other minorities that are scattered all over the country, like Turkish immigrants in many parts of Europa. This is a bit unsatisfactory, but otherwise at least a dozen languages had to be listed for most West European countries.

The same applies to Religions: Not all are listed, but most. Usually the percentages are provided in brackets, but these figures can only be clues. They differ very much from source to source. For some countries you get a total of more than 100%, for others less than 70%. This is no mistake but due to the fact that some people belong to two religious communities, others to none.

5.2 Population data

The number of inhabitants is as up-to-data as possible, its annual growth is the average value for the last decade. The density is not very helpful for some large countries, like Egypt: The Nile delta is one of the most densely populated regions on Earth while the deserts are uninhabited.

The life expectancy is the average value for women and men. The actual figure for women is a few years more, for men a few years less.

The share of illiterates is mostly taken from national authorities of a country and therefore often too little: Even in the industrialized countries of the West, there is a hidden illiteracy of a few percent.

The figure for the proportion of rural and urban population is problematic, either: Different countries have different measuring methods for that one.

5.3 Economic data

The Grand National Product (GNP) is as up-to-date as possible, usually taken from 1995. It is an indicator for the economic strength of a country and composed of the overall production and service in that country, plus money income from abroad, minus money transfer to foreign countries.

The economic growth is the percentage that the GNP has grown within a year. This figure is quite instable and AmigaWorld would always be out of time if using the value for one single year. Therefore the figure is the average value for the past five years, not showing you the tendency an economy has right now but its development in recent years.

The inflation varies even more throughout the years: AmigaWorld offers the average for the past decade.

On the contrary the percentage of unemployment is as up-to-date as possible, usually it is the average value for a whole year, because unemployment is subject to seasonal changes.

The unit of Energy consumption is kilogramme oil units.

6 Maps

The map window has a little menu with the item “Description”, giving you a text about the continent you are looking at.

At the moment there are only ten maps showing the location of a country. Future versions will include maps that flexibly adjust to differnt screen modes and provide more details, like cities, lakes, mountains and so on.

7 Description Texts

You can move in these texts not only with the proportional gadget, but also with the cursor up-/down-keys. You can print out or save the descriptions with a little menu.

The descriptions of the continents and most organizations I have written. Ruud Hoekzema has helped with criticism and ideas, and he has included some more organizations.

But the lion’s share of this work had and still has Bo Arnholm who has written the descriptions for all 258 countries and dependencies. Bo is still improving these texts. If you like to help him: Read the letter that he has written to you. The ASCII-file is included in the same drawer as this documentation.

And please notice that these descriptions are copyright by Bo Arnholm. That means that you have to contact him first if you would like to make any non-private use of them.

8 Legal stuff

8.1 Disclaimer

This software and data is provided “as is”. No representations or warranties are made with respect to the accuracy, reliability, performance, currentness, or operation of this software and data, and all use is at your own risk. The authors do not assume any responsibility or liability whatsoever with respect to your use of this software and data.

8.2 Distribution

Everybody is encouraged to spread the limited demo version of AmigaWorld, as long as he or she does not remove, add or change any of the files.

The “keyfile” obtained by registered users must not be given to other persons. It would be no good idea anyway, because it contains the user’s name. Should you commit this heinous deed then I wish you that you have to sit in front of a computer for the rest of your life!

9 Registration

AmigaWorld has been released as ShareWare. That means that everybody can copy the software, and you can use most of its features without paying. The authors do not get any money unless you become a registered user and send a donation to us.

If you like the software and would like to make its authors happy (and willing to create future versions) then please become a registered user and honor our work!

Please see the file “Registration.asc” in this drawer for more info.

10 Creating a new translation

From the very beginning, AmigaWorld has been designed as a multi-language program. All output messages and all the data is stored in seperate files for each language. AmigaWorld’s “FarawayEyes” font contains the special chars of the Croatian, Czech, Hungarian, Polish and Turkish languages (like ǎ, ş and so on); and other fonts may be used, either. So a translation in all European languages is possible!

Hence it is not very complicated to translate AmigaWorld, but it is a lot of work, as AmigaWorld knows lots of data. If you would like to translate Amiga-World into another language, you have to cope with three files:

• aw.output contains about 200 lines of output messages.

• aw.data contains all the country data.

• aw.org contains the descriptions of international organizations.

Have a look at these files, and if you are still interested in doing a new translation – you are very welcome! But please send a message to me (Wolfgang) first, as somebody else might be working on the same language.

Translating AmigaWorld is mostly a favour to the people that speak your language: Don’t expect to get richt! You will receive your keyfile for free and if you like to, you can be a registration site for your country: Your income might suffice for a good dinner and a nice bottle of wine – once or twice in a year…

11 A small request

Every new version of AmigaWorld is thoroughly tested before its release, and severe bugs have rarely been found. Should such a bug or even a complete crash of the system happen to you then please write me which kind of hardware you use, and what exactly happened. I will be very thankful for hints about wrong data, too! Any comments and suggestions on AmigaWorld are welcome to me,

   Lug@stud-mailer.uni-marburg.de

If you have any comments, suggestions or corrections on the country descriptions, then don’t hesitate to write to Bo Arnholm,

   arnis@tripnet.se

12 A big “Dankeschön”

A real difficulty when saying “Thank you” to more than one person is choosing an order! Who shall be mentioned first and who last? This awful problem has been solved here by using order of appearance, as it is done when listing the actors of a stage play – and in fact, is AmigaWorld anything else but a real good tragedy? If you would like to turn it into a comedy then please become a registered user!

So, I would like to thank everybody who helped with ideas, critizism and their work, especially:

• Alexander Linke who showed me what a good thing Modula-2 is: The AmigaWorld program has been written using this instruction code.

• Bo Arnholm for his translation to Swedish and for all his ideas, and of course for all the wonderful country descriptions that he has written during the last two years.

If you would like to contact him for discussion his texts, or simply for sending him an extra donation: His address is

Bo Arnholm – Södermalmsgatan 13 – 431 69 Mölndal – Schweden

email: arnis@tripnet.se

• Ruud Hoekzema for his Dutch translation and his improving the country data as well as the descriptions of continents and organizations, and for all his good ideas. Thank you also for giving home to one of the AmigaWorld homepages at the great server xs4all.

• Mika Leinonen for his translation to Finnish and for adding some dependencies, and also for providing the first web-address of the AmigaWorld homepage.

• Olivier Bogros und Yves Salingue who together have created the French translation of the program and the documentation, and for maintaining the files in such a brilliant way.

• Dr. Guido Alfani who rendered AmigaWorld into Italian. As a political scientist (and therefore a colleague), he provided lots of tips for the organization descriptions.

• Casper Thomsen for the Danish translation of AmigaWorld and for many good ideas on how to make AmigaWorld even better.

• Bjarne Bjørnsen who translated AmigaWorld into Norwegian, making it possible for all North Europeans to use the program in their mother tongue.

• Dave Travers who added many flags and improved the appearance of most of the old ones.

• Lars Unger who designed and created the web pages of AmigaWorld, with a lot of work and enthusiasm.

I would also like to thank two other authors of non-commercial Amiga software which was very helpful for creating AmigaWorld:

• Jean-Michel Forgeas for his great “Az”. After five years of using, this dinosaur still is my favourite editor, and most AmigaWorld data files have been written with it.

• Thies Wellpott for his excellent “PlotMap”. The outlines of the maps shown by AmigaWorld have originally been created with this software.

A Example

A little example will show you the capabilities of AmigaWorld for comparing data of different countries. It describes how to compare the economic growth of the EU countries to the growth of the East Asian nations.

1. Choose “Organization” from the “Selection” menu. A list of international organizations appears.

2. Click on the European Union. A description of the EU comes up. Read it or not, and close the window.

3. Now choose “Grand National Product / Growth” from the “Sort by” menu. You will see a list of the 15 EU members, beginning with the most booming economies.

4. By activating “Rank” and “Value” from the “Display menu” you can easily compare the figures for the EU countries.

5. Using “Calc Average” from the “Selection” menu will give you the average value.

6. Change the search mode to “include”, and the search for-gadget to “location”.

7. Type in “East Asia”. This will produce a mixed list of the European Union and East Asia, still sorted by economic growth.

8. There is a little trick for using two different colours: Choose “Restore” from the “Selection menu” and “Compare” from the “Display” menu.

Ain’t it “wunderbar”?

B Keyboard commands

Most features of AmigaWorld can be accessed via keyboard, too: The important menus have a shortcut, and the main window offers the following key commands:

Cursor-up/down: Highlights a country and scrolls the entire list when the highlight bar comes to the top or end of the list. Use the cursor-keys in connection with Shift and Alt to move the whole list up and down.

Return: Calls up the country data for the highlighted country.

Space bar: Highlighting off / on again.

Escape: Closes any window except for the main window.

The search gadgets and the selection windows for continents and organizations also provide full keyboard support: Just hit the understroken letter with or without the Shift key.

C AmigaWorld on the WorldWideWeb

Since the end of 1996, AmigaWorld is also represented on the WorldWideWeb! You cannot look at countries’ data or have lists created online – you have got the program for that purpose, anyway. And it’s cheaper to get informed offline.

So, what is the aim and object of our homepage? You can stay up-to-date with program development as new versions of AmigaWorld are announced on the Web and can also be downloaded. What’s best: You can download only those parts that you really need. If you want the latest data files in English, then you can get them and not a single more byte.

Lars Unger has designed and created the pages, and at the moment there are three URLs that have been made available by Mika Leinonen, Ruud Hoekzema and Lars himself:

   • <http://stekt.oulu.fi/~mikal/amigaworld> (in Finland)
   • <http://www.xs4all.nl/~rhoek/amigaworld> (in the Netherlands)
   • <http://www.amigaworld.com/Lars/aw> (in the USA)

If you would like to know what else can be found on the site – well, have a look :-)